

********************************************************************
UARS WEEKLY STATUS REPORT - 07 May 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off on Thursday, May 6. Otherwise, the experiment
successfully gathered its daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1997239. MEPS data files between
1997240 and 1997365 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 24 April 2004 (Orbit 69044, GMT Day 115) through 30
April 2004 (Orbit 69149, GMT Day 121).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD
The beta angle ranged between the angles of 52.1 degrees to 26.5
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 07 May 2004 (DOY 128).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.95 to 0.96 degrees C. Additional
battery performance data for each battery on 23 April and 30 April
2004 are provided in Appendix D.
The clock error ranged from 10.4 to -09.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 03
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT began taking SSA only events on TDW on 23 April 2004. This
change to operations was needed due to the TDRS-8 move to TDW. Since
the SSA service is in higher demand than the MA service, the FOT
receives approximately 50% of the events requested. This effectively
yields the realtime support of 2.5 TDRS instead of 3 TDRS. The FOT
will continue to operate in this mode until the software fix to the
CMS is made to ingest the SMA view periods correctly. Preliminary
estimates project this fix to be released to operations no sooner than
two months from now.
On 29 April 2004, UARS experienced an anomaly in the ATC load. The
wrong iteration of the ATC load was uplinked for this day (DOY 120),
which caused the ATC to execute improperly. The results of this error
were a missed realtime event (120/2000-2024 Z), 46 missed stored
commands (which included a HALOE record session) and all SSPP data
from 120/0000-2200 Z. The error was noticed when data was not
acquired on the 120/2000 event, and the mistake was corrected on the
following 120/2108 event. The FOT apologizes to the instruments teams
whose data was lost, and is reviewing actions taken and online
procedures to insure this does not happen again.
MEETINGS
The FOT attended a meeting with the HRDI team and FDF to discuss the
attitude error seen in the HRDI data. A working group meeting has
been scheduled in two weeks. In the meantime, FDF has been given some
requested attitude data to review in preparation for this meeting.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 07 May 2004 (DOY 128). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 April 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
69048 115/0558 04-124 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69140 121/0940 04-125 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 April 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 April 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
69048 115/0558 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69140 121/0940 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 April 2004
23 April 2004 - GMT Day 114 Beta = 53.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-089.6 +3.87/+2.60 +22.8/-14.4 25.8 1.05 15.5
3 +00.0/-123.2 +2.91/+1.65 +22.4/-13.6 25.8 1.05 15.1
30 April 2004 - GMT Day 121 Beta = 26.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-179.2 +5.46/+4.50 +28.4/-14.0 24.6 1.04 20.4
3 +22.4/-134.4 +4.50/+3.55 +27.6/-14.4 24.6 1.03 20.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 April 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-021 FOT 69060 116/0050 30824 Telemetry Dropouts
04-022 FOT 69132 120/2000 35744 Negative Acquisition
04-023 FOT 69134 120/2331 35746 Data Drop Out
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-020 FOT 68977 110/1337 N/A Eclipse False Night Indicator
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-019 FOT See List Apr 04 30832 Generic Late Acquisition
04-021 FOT 69060 116/0050 30824 Telemetry Dropouts
04-022 FOT 69132 120/2000 35744 Negative Acquisition
04-023 FOT 69134 120/2331 35746 Data Drop Out